Your car burns 2 1/2 at of oil in 5000 miles how much oil will it burn in 7200 miles

Found 2 solutions by amalm06, stanbon:
Answer by amalm06(224)

Ratio problem:

(2.5 oil/5000 mi)*(7200 mi)=3.6 oil (Answer)

Answer by stanbon(75887)

Your car burns 2 1/2 at of oil in 5000 miles how much oil
will it burn in 7200 miles
----
Use a proportion::
x/7200 = 2.5/5000
---
x = 7200*25/50000
x = 72*25/500
x = 72*1/20
x = 36/10
x = 3.6 gallons
-------------
Cheers
Stan H.